# Spectral Bottleneck in Deep Neural Networks: Noise is All You Need

### v1

@@ 3 WINNER: Weight Initialization with Noise for NEural Representations
In Sec. 2.2, we have emphasized that the root cause of the spectral bottleneck (Fig. 4) is the mismatch of spectral energy between the target signal and network activations at initialization. We address this issue using Proposition 1 to alter the Fourier representation of the network’s activations and outputs at initialization. To this end, we introduce WINNER, a weight perturbation scheme in which a Gaussian noise is added to the uniformly initialized weights that exist between the inputs and the second hidden layer. So, the weights immediately upstream of first and second hidden layers are perturbed as,
@@
$\text{SIREN}^{2}:f(\mathbf{x};\theta)=\mathbf{W}^{(L)}\mathbf{h}^{(L-1)}+\mathbf{b}^{(L)},\quad\mathbf{h}^{(l)}=\begin{cases}\mathbf{x},&l=0\text{ (inputs)}\\ \phi^{\text{sin}}\big{(}(\boldsymbol{\mathbf{W}}^{(l)}+\eta_{jk}^{(l)})\mathbf{h}^{(l-1)}+\mathbf{b}^{(l)}\big{)},&l=1,2,\\ \phi^{\text{sin}}\big{(}\mathbf{W}^{(l)}\mathbf{h}^{(l-1)}+\mathbf{b}^{(l)}\big{)},&l=3,\dots,L-1,\end{cases}$ (9)
Although the weight perturbations are added only up to the second hidden layer, their effect propagates downstream all the way to the outputs. This is shown empirically in Sec. 4 and in Supplementary Material Sec. A for the full network. The goal of the proposed noise addition scheme is to enhance the functional sensitivity between the outputs and network parameters necessary to allow the parameter updates required for fitting high-frequency modes.
Figure 5: Distributions of the dot product between inputs $X_{ij}\sim\text{Arcsin}(-1,1)$ and weights initialized under two initialization schemes: (a) standard uniform weights $W_{jk}=\sim\mathcal{U}\left(-\omega_{0}^{-1}\sqrt{6/d},\omega_{0}^{-1}\sqrt{6/d}\right)$, and (b) WINNER, in which uniform weights are perturbed with Gaussian noise, $W_{jk}+\eta_{jk}$. The noise addition increases the standard deviation of the dot product from $1$ to $\sqrt{1+\frac{ds^{2}}{2}}$, where $d$ is the input dimension (fan_in). This increase closely matches the analytically predicted value (Theorem 3.1) shown by the dashed black line.
@@
$\mathrm{Var}[X_{ij}\eta_{jk}]=\mathbb{E}[(X_{ij}\eta_{jk})^{2}]-(\mathbb{E}[X_{ij}\eta_{jk}])^{2}=\mathbb{E}[X_{ij}^{2}]\cdot\mathbb{E}[\eta_{jk}^{2}]-0=\frac{1}{2}\cdot\left(\frac{s^{2}}{\omega_{0}^{2}}\right).$
@@
3.1 Target-aware Specification of Noise Scales $s_{0}$ and $s_{1}$ for SIREN2
@@
The performance of INRs is highly sensitive to the spectral content of the target signal (Figs. 1, 2, and 4). Therefore, an effective weight initialization is one that is target-aware and accounts for the spectral profile of the target. To quantify the distribution profile of different frequencies of the target, the spectral centroid $\psi$ is defined as the normalized average frequency of the target’s power spectrum, computed as
$\psi=2\times\frac{\sum k|\hat{y}(k)|}{\sum_{|}\hat{y}(k)|},$ (10)
where $\hat{y}(f_{k})$ is the Fourier transform of the target signal evaluated at frequency bin $k$. The factor $2$ normalizes $\psi$ to the range $[0,1]$. Using $\psi$, the noise scales $s_{0}$ and $s_{1}$ in Eqn. 8 are defined as
$s_{0}=s_{0}^{\text{max}}\!\left(1-e^{\,a\frac{\psi}{C}}\right),\quad s_{1}=b\left(\frac{\psi}{C}\right),$ (11)
where $C$ denotes the number of channels. The hyperparameters $s_{0}^{\text{max}},a,b]$ are chosen as $[3500,5,3]$ for audio fitting and $[50,5,0.4]$ for image fitting.
4 Spectral Properties of SIREN2
4.1 Neural Tangent Kernel at Initialization
@@
The spectral characteristics of the NTK at initialization are examined with an aim to understand the frequency support of SIREN2 in comparison to SIREN. As shown in Fig. 7, SIREN exhibits rapid eigenvalue decay and allocates higher cumulative spectral energy $\mathcal{S}(k)$ to the low-frequencies. This rapid collapse of energy occurs due to the alignment of dominant NTK eigenvectors with smooth, low-frequency functions, restricting expressivity in tasks requiring fine-scale resolution such as audio representation. The proposed WINNER used in SIREN2 provides significant improvement, featuring a tunable $\mathcal{S}(k)$ profile controlled by weight perturbation scales $s_{0}$ and $s_{1}$, analogous to the Fourier scales used in random Fourier embeddings rahimi2007random ; tancik2020fourier . By contrast, SIREN2 captures high-frequencies and positional information directly through its sinusoidal activations, avoiding the quadratic increase in parameters required by random Fourier or positional embeddings to represent similar frequency content. As demonstrated in Fig. 7, SIREN2 can be configured to exhibit slower eigenvalue decay with higher cumulative spectral energy $\mathcal{S}(k)$ distributed across the entire spectrum with parameters $s_{0}$ and $s_{1}$. The critical factor for a successful implicit representation is the appropriate selection of $s_{0}$ and $s_{1}$ (Sec. 3.1).
@@
Figure 7: Controlling the NTK spectra via noise scales $s_{0}$ and $s_{1}$ of WINNER. a Normalized NTK kernels for SIREN and SIREN2 with $s_{0},s_{1}]=[10,0.5]$, showing reduced off-diagonal correlations. b NTK eigenvalue decay profile (left) and eigenvalue-weighted FFT magnitude spectra - $\mathcal{S}(k)$ (right) for varying noise scales, demonstrating that SIREN2 broadens frequency support. All networks shown here employ four hidden layers with 256 features, use $\omega_{0}=30$, and are evaluated on $2^{10}$ uniformly sampled inputs in [1,1][-1,1].
4.2 Activation Spectra
@@
Figure 8 compares the distributions and power spectral densities (PSDs) of network inputs and the pre- and post-activation values in the first two hidden layers for SIREN and SIREN2. For SIREN, pre-activation distributions are approximately Gaussian, while post-activations follow the $\text{Arcsin}(-1,1)$ law, consistent across layers. SIREN2 preserves these distributional structure, albeit with broader pre-activation spreads in layers 1 and 2. Distributions for subsequent layers (not shown) remain identical to that of SIREN; full-layer distributions are reported in the Supplementary (Sec. A).
@@
5 Experiments
We evaluate the performance of SIREN2 (SIREN initialized with WINNER) against several state-of-the-art INR architectures from the literature, including the baseline SIREN sitzmann2020implicit , Gauss ramasinghe2022beyond (2022), WIRE saragadam2023wire (2023), and FINER liu2024finer (2024), in reconstructing a variety of challenging audio signals. Following the recommendation in sitzmann2020implicit , all architectures, except SIREN2, use a scaling factor of 100 for the first-layer activation periodicity in the audio fitting experiments.
5.1 1D Audio Fitting
To ensure consistency across experiments, all audio signals are fixed at a length of 150,000 samples, making the setup invariant to sampling rate. Each network is trained for 30,000 epochs, and every experiment is conducted over five random trials to compute the mean peak PSNR and standard deviation. The selected audio signals span a range of spectral characteristics, including high-frequency-dominant, low-frequency-dominant, and broadband signals.
Table 1 shows that the proposed SIREN2 delivers consistently higher reconstruction accuracy than existing INR architectures across diverse audio signals, establishing new state-of-the-art results. These gains are especially notable for signals with strong high-frequency content, where other methods exhibit substantial residual errors. While the weight perturbation scheme amplifies high-frequency energy (Fig. 5), it slightly suppresses low frequencies, as seen in the reconstruction of bach.wav. This indicates the need for a more robust choice of $s_{0}$ and $s_{1}$ for such low-frequency dominant signals. As illustrated in Fig. 9, SIREN2 enables high-fidelity audio reconstructions, achieving PSNR values above $60$ dB with parameter count comparable to that of number of samples in the signal.
Table 1: Audio fitting. Mean and standard deviation of PSNR values for different architectures on audio signal reconstruction. The proposed SIREN2 achieves state-of-the-art performance. Results are color coded as best, second best, and third best reconstructions. The network width is chosen so that the total parameter count is approximately equal to the signal length.
@@ 5 Experiments
Figure 9: High-fidelity audio reconstruction with SIREN2. Each column corresponds to an audio signal: tetris.wav, sparking.wav, and shattered.wav. Top row: ground truth spectrograms. Second row: PSNR histories for different network architectures. Remaining rows show the spectrogram errors for reconstructions by each model. SOS-SIREN stands out by exhibiting near-zero reconstruction noise.
@@
5.2 2D Image Fitting
We evaluate SIREN2 on 2D image fitting tasks, $f(\mathbf{x};\theta):\mathbb{R}^{2}\mapsto\mathbb{R}^{d}$, with $d=1$ for grayscale and $d=3$ for RGB images. The experiments cover a diverse set of images, including natural images from the Kodak kodak_photocd dataset, challenging texture images from two DTD cimpoi14describing classes (braided and woven), and synthetic high-frequency patterns. Table 2 shows that SIREN2 consistently outperforms the original SIREN across all cases, with PSNR improvements ranging from 7% (2D-Riemann) to 69% (noise1.png), and especially strong gains for images with high-frequency content and small pixel count (in the over-paramaterized regime). For RGB image reconstruction, SIREN2 provides only marginal gains over SIREN, while FINER achieves the best performance. Figure 10 visualizes these improvements using FFT error maps, highlighting that SIREN2 achieves lower fitting error in high-frequency regions. The frequency-domain analysis further confirms that SIREN2 preserves fine details and sharp transitions more accurately, as indicated by the reduced magnitude errors (darker regions).
Table 2: 2D image fitting. PSNR($\uparrow$) in dB across images and datasets for different INR architectures. SIREN2 consistently surpasses SIREN, with percentage gains (in parentheses) reflecting improvements achieved purely through initialization. Results are color coded as best, second best, and third best reconstructions.
@@
Figure 10: Accuracy improvements in Fourier space with SIREN2 for image fitting. Reconstruction performance of SIREN and SIREN2 models for various images. The top two rows show the ground truth images and their corresponding FFT magnitudes. Subsequent rows depict the FFT error maps and peak PSNR values achieved by different models. SIREN2 consistently produce lower FFT errors and higher PSNRs.
@@
Reproducibility details. All networks follow the same settings as the audio fitting experiments, with no scaling applied to the first layer $\omega_{0}$. The ReLU+PE presented in Table 2 incorporates positional encoding with 256 embeddings. The positional encodings use a logarithmic frequency spectrum, with frequencies ranging from $2^{0}$ to $2^{\text{n}-1}$, with n=7 (number of frequencies).
5.3 Image denoising
Figure 11: 2D denoising. Reconstruction of a noisy density field from a 2D Riemann problem kurganov2002solution ; chandravamsi2024high . The ground truth corrupted with Gaussian noise (top middle) and reconstructions using different networks are shown. SIREN2 achieves comparatively the best fidelity. Denoising accuracy is evaluated using PSNR and structural similarity index (SSIM) with respect to the ground truth.

### v2

@@ IV WINNER: Weight Initialization with Noise for NEural Representations
In Sec. III-B, we have emphasized that the root cause of the spectral bottleneck (Fig. 3) is the mismatch of spectral energy between the target signal and network activations at initialization. We address this issue using Proposition 1 to alter the Fourier representation of the network’s activations and outputs at initialization. To this end, we introduce WINNER, a weight perturbation scheme in which a Gaussian noise is added to the uniformly initialized weights that exist between the inputs and the second hidden layer. So, the weights immediately upstream of first and second hidden layers are perturbed as,
@@
$\begin{gathered}\text{SIREN}^{2}:\;f(\mathbf{x};\theta)=\mathbf{W}^{(L)}\mathbf{h}^{(L-1)}+\mathbf{b}^{(L)},\\[4.0pt] \mathbf{h}^{(l)}=\begin{cases}\mathbf{x},&l=0\text{ (inputs)},\\[3.0pt] \phi^{\text{sin}}\!\big((\mathbf{W}^{(l)}+\eta_{jk}^{(l)})\mathbf{h}^{(l-1)}+\mathbf{b}^{(l)}\big),&l=1,2,\\[3.0pt] \phi^{\text{sin}}\!\big(\mathbf{W}^{(l)}\mathbf{h}^{(l-1)}+\mathbf{b}^{(l)}\big),&l=3,\dots,L-1.\end{cases}\end{gathered}$ (9)
Although the weight perturbations are added only up to the second hidden layer, their effect propagates downstream all the way to the outputs. This is shown empirically in Sec. V and in Supplementary Material Sec. LABEL:suppl:fulldists for the full network. The goal of the proposed noise addition scheme is to enhance the functional sensitivity between the outputs and network parameters necessary to allow the parameter updates required for fitting high-frequency modes.
Figure 5: Distributions of the dot product between inputs $X_{ij}\sim\text{Arcsin}(-1,1)$ and weights initialized under two initialization schemes: (a) standard uniform weights $W_{jk}=\sim\mathcal{U}\left(-\omega_{0}^{-1}\sqrt{6/d},\omega_{0}^{-1}\sqrt{6/d}\right)$, and (b) WINNER, in which uniform weights are perturbed with Gaussian noise, $W_{jk}+\eta_{jk}$. The noise addition increases the standard deviation of the dot product from $1$ to $\sqrt{1+\frac{ds^{2}}{2}}$, where $d$ is the input dimension (fan_in). This increase closely matches the analytically predicted value (Theorem IV.1) shown by the dashed black line.
@@
$\mathrm{Var}[X_{ij}\eta_{jk}]$ $=\mathbb{E}[(X_{ij}\eta_{jk})^{2}]-(\mathbb{E}[X_{ij}\eta_{jk}])^{2}$
$=\mathbb{E}[X_{ij}^{2}]\cdot\mathbb{E}[\eta_{jk}^{2}]-0=\frac{1}{2}\cdot\left(\frac{s^{2}}{\omega_{0}^{2}}\right).$
@@
IV-A Target-aware Specification of Noise Scales $s_{0}$ and $s_{1}$
As shown in Figs. 1, 2, and 3, the performance of INRs is sensitive to the spectral content of the target signal. Therefore, an effective weight initialization is one that is target-aware and accounts for the spectral profile of the target. To quantify the weighted contribution of different frequencies to the target, a spectral centroid $\psi$ is defined as the normalized average frequency of the target’s power spectrum, computed as
$\psi=2\times\frac{\sum_{k}k|\hat{y}(k)|}{\sum_{k}|\hat{y}(k)|},$ (10)
where $\hat{y}(f_{k})$ is the Fourier transform of the target evaluated at frequency bin $k$. The factor $2$ normalizes $\psi$ to the range $[0,1]$. Using $\psi$, the noise scales $s_{0}$ and $s_{1}$ in Eqn. 8 are empirically determined as
$s_{0}=s_{0}^{\text{max}}\!\left(1-e^{-a\frac{\psi}{C}}\right),\quad s_{1}=s_{1}^{\text{max}}\left(\frac{\psi}{C}\right),$ (11)
where $C$ denotes the number of channels (e.g., $C=3$ for RGB images). The hyperparameters $s_{0}^{\text{max}},a,s_{1}^{\text{max}}]$ are set as $\left[\frac{3500}{70^{N_{d}-1}},5,\frac{4}{10^{N_{d}-1}}\right]$, with $N_{d}$ representing the input dimension of the network. Eqn. 11 is valid for all modalities including audio, images, video, and 3D data.
The functional form of $s_{0}$ in Eqn. 11 is empirically determined through multiple experiments. An initial linear formulation was tested, but experiments relating the spectral centroid to optimal $s_{0}$ values revealed a trend better captured by an exponential curve rather than a linear one (see Supplementary Fig. S1). The linear dependence was retained for $s_{1}$. Both expressions are therefore ad-hoc and derived from multiple empirical evaluations; further work is required to develop a principled, target-aware formulation. Additional plots illustrating the sensitivity of noise scales to fitting accuracy for various audio clips are included in the supplementary material. As also detailed in sitzmann2020implicit, the larger input scaling factor or the larger noise scales used for audio signals arise from their high sampling rates and higher frequency content compared to visual data.
@@ IV-A Target-aware Specification of Noise Scales $s_{0}$ and $s_{1}$
SIREN2 vs. FINER++: The current work compares against FINER/FINER++ liu2024finer; zhu2024finerplusplus, which also aim to improve high-frequency representation but do so by modifying the bias initialization and the activation function. However, altering biases does not mathematically guarantee Gaussian-distributed pre-activations, and therefore cannot inherit the properties of Gaussian initialization that stabilize optimization and control signal propagation glorot2010understanding. Another key difference is that FINER++ does not modify the weights connecting the first and second hidden layers, which limits its ability to adjust the initial spectral profile of pre-activations downstream of the first hidden-layer.
V Spectral Properties of SIREN2
V-A Neural Tangent Kernel at Initialization
@@
The spectral characteristics of the NTK at initialization are examined with an aim to understand the frequency support of SIREN2 in comparison to SIREN. As shown in Fig. 7, SIREN exhibits rapid eigenvalue decay and allocates higher cumulative spectral energy $\mathcal{S}(k)$ to the low-frequencies. This rapid collapse of energy occurs due to the alignment of dominant NTK eigenvectors with smooth, low-frequency functions, restricting expressivity in tasks requiring fine-scale resolution such as audio representation. The proposed WINNER used in SIREN2 provides significant improvement, featuring a tunable $\mathcal{S}(k)$ profile controlled by weight perturbation scales $s_{0}$ and $s_{1}$, analogous to the Fourier scales used in random Fourier embeddings rahimi2007random; tancik2020fourier. By contrast, SIREN2 captures high-frequencies and positional information directly through its sinusoidal activations, avoiding the quadratic increase in parameters required by random Fourier or positional embeddings to represent similar frequency content. As demonstrated in Fig. 7, SIREN2 can be configured to exhibit slower eigenvalue decay with higher cumulative spectral energy $\mathcal{S}(k)$ distributed across the entire spectrum with parameters $s_{0}$ and $s_{1}$. The critical factor for a successful implicit representation is the appropriate selection of $s_{0}$ and $s_{1}$ (Sec. IV-A).
@@
Figure 7: Controlling the NTK spectra via noise scales $s_{0}$ and $s_{1}$ of WINNER. (a) Normalized NTK kernels for SIREN and SIREN2 with $s_{0},s_{1}]=[10,0.5]$, showing reduced off-diagonal correlations. (b) NTK eigenvalue decay profile (left) and eigenvalue-weighted FFT magnitude spectra - $\mathcal{S}(k)$ (right) for varying noise scales, demonstrating that SIREN2 broadens frequency support. All networks shown here employ four hidden layers with 256 features, use $\omega_{0}=30$, and are evaluated on $2^{10}$ uniformly sampled inputs in [1,1][-1,1].
V-B Activation Spectra
@@
Figure 8 compares the distributions and power spectral densities (PSDs) of network inputs and the pre- and post-activation values in the first two hidden layers for SIREN and SIREN2. For SIREN, pre-activation distributions are approximately Gaussian, while post-activations follow the $\text{Arcsin}(-1,1)$ law, consistent across layers. SIREN2 preserves these distributional structure, albeit with broader pre-activation spreads in layers 1 and 2. Distributions for subsequent layers (not shown) remain identical to that of SIREN; full-layer distributions are reported in the Supplementary (Sec. LABEL:suppl:fulldists).
@@
VI Experiments
We evaluate the performance of SIREN2 (SIREN initialized with WINNER) against several state-of-the-art INR architectures from the literature, including the baseline SIREN sitzmann2020implicit, Gauss ramasinghe2022beyond (2022), WIRE saragadam2023wire (2023), and FINER liu2024finer (2024), in reconstructing a variety of challenging audio signals. Following the recommendation in sitzmann2020implicit, all architectures, except SIREN2, use a scaling factor of 100 for the first-layer activation periodicity for the audio fitting experiments.
VI-A 1D Audio Fitting
To ensure consistency across experiments, all audio signals are fixed at a length of 150,000 samples, making the experimental setup invariant to sampling rate. Each network is trained for 30,000 epochs, and every experiment is conducted over five random trials to compute the mean peak PSNR and standard deviation. The selected audio signals span a range of spectral characteristics, including high-frequency-dominant, low-frequency-dominant, and broadband signals.
Table I shows that the proposed SIREN2 delivers consistently higher reconstruction accuracy than existing INR architectures across diverse audio signals, establishing new state-of-the-art results. These gains are especially notable for signals with strong high-frequency content, where other methods exhibit substantial residual errors. While the weight perturbation scheme amplifies high-frequency energy (Fig. 5), it slightly suppresses low frequencies, as seen in the reconstruction of bach.wav. This indicates the need for a more robust choice of $s_{0}$ and $s_{1}$ for such low-frequency dominant signals. As illustrated in Fig. 9, SIREN2 enables high-fidelity audio reconstructions, achieving PSNR values above $60$ dB with parameter count comparable to that of number of samples in the signal.
TABLE I: Audio fitting. Mean and standard deviation of PSNR values for different architectures on audio signal reconstruction. The proposed SIREN2 achieves state-of-the-art performance. Results are color coded as best, second best, and third best reconstructions. The network width is chosen so that the total parameter count is approximately equal to the signal length.
@@ VI Experiments
Figure 9: High-fidelity audio reconstruction with SIREN2. Each column corresponds to an audio signal: tetris.wav, sparking.wav, and shattered.wav. Top row: ground truth spectrograms. Second row: Averaged PSNR histories from different network architectures with the region of standard deviation shaded. Remaining rows show the spectrogram errors for reconstructions by each model. SIREN2 stands out by exhibiting near-zero reconstruction noise.
@@
VI-B 2D Image Fitting
We evaluate SIREN2 on 2D image fitting tasks, $f(\mathbf{x};\theta):\mathbb{R}^{2}\mapsto\mathbb{R}^{d}$, with $d=1$ for grayscale and $d=3$ for RGB images. The experiments cover a diverse set of images, including natural images from the Kodak kodak_photocd dataset, challenging texture images from two DTD cimpoi14describing classes (braided and woven), and synthetic high-frequency patterns. Table II shows that SIREN2 consistently outperforms the original SIREN across all cases, with PSNR improvements ranging from 7% (2D-Riemann) to 69% (noise.png), and especially strong gains for images with high-frequency content and small pixel count (in the over-paramaterized regime). For RGB image reconstruction, SIREN2 provides only marginal gains over SIREN, while FINER achieves the best performance. Figure 10 visualizes these improvements using FFT error maps, highlighting that SIREN2 achieves lower fitting error in high-frequency regions. The frequency-domain analysis further confirms that SIREN2 preserves fine details and sharp transitions more accurately, as indicated by the reduced magnitude errors (darker regions).
TABLE II: 2D image fitting. PSNR($\uparrow$) in dB across images and datasets for different INR architectures. SIREN2 consistently surpasses SIREN, with percentage gains (in parentheses) reflecting improvements achieved purely through initialization. Results are color coded as best, second best, and third best reconstructions.
@@
Figure 10: Accuracy improvements in Fourier space with SIREN2 for image fitting. Reconstruction performance of SIREN and SIREN2 models for various images. The top two rows show the ground truth images and their corresponding FFT magnitudes ($|\mathcal{F}(f_{gt})|$). Subsequent rows depict the FFT error maps and peak PSNR values achieved by different models. SIREN2 consistently produce lower FFT errors and higher PSNRs.
@@
Reproducibility details. All networks follow the same settings as the audio fitting experiments, with no scaling applied to the first layer $\omega_{0}$. The ReLU+PE presented in Table II incorporates positional encoding with 256 embeddings. The positional encodings use a logarithmic frequency spectrum, with frequencies ranging from $2^{0}$ to $2^{\text{n}-1}$, with n=7 (number of frequencies).
VI-C Image Denoising
Figure 11: Image denoising. Comparison of reconstructions for noisy 2D fields and natural images. Rows 1–2: reconstruction of a noisy density field obtained from a computational fluid dynamics simulation of a 2D Riemann problem kurganov2002solution; chandravamsi2024high. Rows 3–4: reconstruction of a sparrow image and a nebula image. SIREN2 yields the highest fidelity among the compared methods, with quantitative results summarized in Table III.